chris has three pieces of string. The total length is 111cm. The sum of the two larger pieces is 92cm. The middle piece is twice as long as the shortest piece. What are the length of the three pieces?
Answer by Alan3354(69443)

Lengths = a, b & c
a + b + c = 111
a + b = 92
--> c = 19 cm
----
b = 2a
a + 2a = 92
a = 92/3 cm
b = 184/3 cm